Problem 22.34 A cube has sides of length L 30.310Both parts please
Answer #1

electric flux is given by

flux = E.A

flux = E*A*cos(theta)

theta angle between E and A.

E = - 4.31 i + 2.71 k

surface area of S1 = 0.31^2 (-i)

flux through S1 = - 4.31*(- 0.31^2) = .414 N.m/C

surface area of S2 = 0.31^2 k

flux through S2 = 2.71* 0.31^2 = 0.26 N.m/C

surface area of S3 = 0.31^2 j

flux through S3 = 0* 0.31^2 = 0 N.m/C

surface area of S4 = 0.31^2 (-k)

flux through S4 = - 2.71* 0.31^2 = - 0.26 N.m/C

surface area of S5 = 0.31^2 i

flux through S5 = - 4.31* 0.31^2 = - 0.414 N.m/C

surface area of S6 = 0.31^2 (-j)

flux through S6 = 0*0.31^2 = 0 N.m/C

B)

total electric flux = Q/eo

total electric flux = 0.414 + 0.26 + 0 - 0.414 - 0.26 + 0 = 0

Q = 0